solve for x
(x-2)/(x-1)=5/x, x#0, x#1

Answer by stanbon(75887)

solve for x
(x-2)/(x-1)=5/x, x#0, x#1
----
Cross-multiply to get:
x^2-2x = 5x-5
---
x^2-7x+5 = 0
----
x = [7 +-sqrt(49-4*5)]/2
----
x = [7 +- sqrt(29)]/2
-----
Cheers,
Stan H.
===========